A tank with 43.2 kg of water is leaking at a rate of 0.0135 kg/s. How many hours will it take until the tank is completely empty?
1
Expert's answer
2019-09-04T03:22:29-0400

t=(m/v) /3600=(43,2/0,0135)/3600=0,89h